### v1

@@ 5.4 Semantic Robustness and Failure Characterization
Semantic Error Taxonomy and Cross-Model Distribution.
To understand why many syntactically valid patches nonetheless fail semantically, we analyze failure cases using a taxonomy of five high-level semantic error types (Table 4). This categorization enables structured evaluation across models and highlights recurring patterns of semantic fragility.
Error Class Description
Requirement Mismatch Misalignment with the intended instruction
Table 4: Semantic error classes and their typical manifestations.
Figure 11: Distribution of semantic error types across models (log-scaled y-axis). Component omission is the most common error overall. GPT-4o and o3-mini frequently produce incomplete or mathematically incorrect edits, while Claude Sonnet 3.7 (thinking) exhibits the lowest error incidence.
Figure 11 presents the frequency of each error type as a proportion of failure cases per model. The most pervasive issue is component omission, accounting for a significant share of failures across all models. This indicates a widespread tendency to under-produce or neglect structurally necessary elements—such as required definitions, lemmas, or proofs—even when the instruction clearly specifies them. Beyond omission, GPT-4o and o3-mini suffer from notably high rates of incompleteness and mathematical errors. These failures suggest limited planning ability when tasked with multi-step edits, and difficulty in maintaining formal soundness under structural pressure. By contrast, Claude Sonnet 3.7 (thinking) demonstrates a markedly lower error profile across all categories.
Summary.
Our findings indicate that semantic failures in LLM-generated patches are not random but structurally patterned. The majority stem from incomplete or underspecified edits, often due to scale underestimation or partial instruction comprehension. While fuzzy repair can address syntactic alignment, it cannot compensate for missing logic or misaligned intent. Addressing these issues will require models that plan multi-step edits, verify task satisfaction, and iteratively refine their output—an agenda we return to in our discussion of future directions.
6 Conclusion
This work introduces the paradigm of Automated Proof Engineering (APE) and presents APE-Bench I, the first large-scale benchmark grounded in real-world proof development workflows from Mathlib4. By reframing formal mathematics as instruction-driven code editing, and combining Lean compilation with LLM-based semantic evaluation, APE-Bench I offers a rigorous and scalable framework for evaluating the structural and semantic capabilities of LLMs in realistic proof engineering tasks. Our empirical results reveal significant performance gaps among state-of-the-art models, particularly under structurally complex scenarios, and underscore the need for future systems that can plan, verify, and refine edits across evolving formal libraries. This benchmark establishes a strong foundation for developing next-generation autonomous agents capable of practical and scalable formal proof maintenance.
Appendix A DiffRepair: Fault-Tolerant Patch Recovery for LLM Outputs
A.1 Motivation and Overview
Model-generated patches in unified diff format frequently fail to apply cleanly to real-world codebases due to misaligned line anchors, outdated contextual information, or excessive reliance on idealized formatting. This undermines the practical viability of LLM-generated edits in structured formal environments like Lean. To address this, we develop DiffRepair, a two-stage recovery system that transforms noisy model diffs into clean, structure-preserving, and context-valid patches that can be compiled and semantically evaluated.
DiffRepair adopts a tolerant matching and reconstruction strategy. Rather than enforcing exact line-based alignment, it reconstructs the intent behind the edit, anchors it in the correct structural region, and regenerates a diff conforming to unified format constraints. This allows it to serve as a bridge between neural code generation and downstream compilation pipelines.
A.2 Algorithmic Workflow
1. Intent Localization.
Given a noisy diff, DiffRepair first identifies where in the original Lean file the intended modification should take place. It uses a multi-granularity fuzzy matching strategy based on contextual anchors surrounding each diff hunk. Anchors are matched using:
  • Exact or prefix-based line similarity.
  • Normalized token-level matching with whitespace and casing tolerance.
  • Structural proximity, emphasizing declaration-aligned contexts (e.g., def, theorem blocks).
Matches are scored using a dynamic programming model optimizing for structural alignment and modification preservation, ensuring minimal distortion to the intended patch intent.
2. Patch Reconstruction.
Once the target region is localized, DiffRepair reconstructs a clean diff by:
  • Re-aligning added and deleted lines to structurally valid positions.
  • Augmenting missing context to satisfy unified diff constraints.
  • Resolving line number offsets and conflicting hunk overlaps.
  • Merging multiple hunks into a minimal, coherent patch that preserves original semantics.
Reconstructed patches are guaranteed to apply via standard tools such as patch or git apply, enabling seamless integration into downstream pipelines.
A.3 Design Features
DiffRepair is designed to satisfy the following desiderata:
  • Robustness to noisy generation: Handles line number inaccuracies, formatting inconsistencies, and incomplete context.
  • Structure preservation: Ensures generated diffs align with Lean’s structural boundaries and avoid cross-declaration contamination.
  • Automation-ready: Requires no human-in-the-loop intervention and supports batch repair at scale.
  • Efficiency: Performs fast multi-hunk matching and reconstruction with low overhead.

### v2

@@ C.3 DiffRepair: Patch Normalization System
As shown in Table 6, raw apply success rates remain low for all models, ranging from 12% to 16%, indicating poor structural precision in the initial output. Exact match repair offers moderate improvement for DeepSeek models, but fails to resolve structural errors in o3-mini or GPT-4o outputs. Fuzzy match repair significantly boosts success across the board, particularly for DeepSeek R1, Doubao, and Claude Sonnet 3.7. These results highlight the necessity of structural repair mechanisms for enabling meaningful evaluation and practical usability.
Appendix D Illustrative Examples
Source code:
Copyright (c) 2017 Mario Carneiro. All rights reserved.
Released under Apache 2.0 license as described in the file LICENSE.
Authors: Mario Carneiro
import Mathlib.Data.Fintype.EquivFin
import Mathlib.Data.Finset.Option
# fintype instances for option
assert_not_exists MonoidWithZero MulAction
open Function
open Nat
universe u v
variable β : Type*}
open Finset Function
instance : Type*} [Fintype α] : Fintype (Option α) :=
Finset.insertNone univ, fun a => by simp
theorem univ_option : Type*) [Fintype α] : (univ : Finset (Option α)) = insertNone univ :=
rfl
@[simp]
theorem Fintype.card_option : Type*} [Fintype α] :
Fintype.card (Option α) = Fintype.card α + 1 :=
(Finset.card_cons (by simp)).trans <| congr_arg _ (card_map _) rfl
/– If Option α‘ is a Fintype then so is ‘α‘ -/
def fintypeOfOption : Type*} [Fintype (Option α)] : Fintype α :=
Finset.eraseNone (Fintype.elems := Option α)), fun x =>
mem_eraseNone.mpr (Fintype.complete (some x))⟩
/– A type is a Fintype if its successor (using Option‘) is a Fintype‘. -/
def fintypeOfOptionEquiv [Fintype α] (f : α Option β) : Fintype β :=
haveI := Fintype.ofEquiv _ f
fintypeOfOption
namespace Fintype
/– A recursor principle for finite types, analogous to Nat.rec‘. It effectively says
that every Fintype is either Empty or Option α‘, up to an Equiv‘. -/
def truncRecEmptyOption {P : Type u Sort v} (of_equiv : β}, α β P α P β)
(h_empty : P PEmpty) (h_option : {α} [Fintype α] [DecidableEq α], P α P (Option α))
: Type u) [Fintype α] [DecidableEq α] : Trunc (P α) := by
suffices n : ℕ, Trunc (P (ULift <| Fin n)) by
apply Trunc.bind (this (Fintype.card α))
intro h
apply Trunc.map _ (Fintype.truncEquivFin α)
intro e
exact of_equiv (Equiv.ulift.trans e.symm) h
apply ind where
Porting note: do a manual recursion, instead of induction tactic,
to ensure the result is computable
/– Internal induction hypothesis -/
ind : n : ℕ, Trunc (P (ULift <| Fin n))
| Nat.zero => by
have : card PEmpty = card (ULift (Fin 0)) := by simp only [card_fin, card_pempty,
card_ulift]
apply Trunc.bind (truncEquivOfCardEq this)
intro e
apply Trunc.mk
exact of_equiv e h_empty
| Nat.succ n => by
have : card (Option (ULift (Fin n))) = card (ULift (Fin n.succ)) := by
simp only [card_fin, card_option, card_ulift]
apply Trunc.bind (truncEquivOfCardEq this)
intro e
apply Trunc.map _ (ind n)
intro ih
exact of_equiv e (h_option ih)
Porting note: due to instance inference issues in SetTheory.Cardinal.Basic
I had to explicitly name h_fintype in order to access it manually.
was ‘[Fintype α]‘
/– An induction principle for finite types, analogous to Nat.rec‘. It effectively says
that every Fintype is either Empty or Option α‘, up to an Equiv‘. -/
@[elab_as_elim]
theorem induction_empty_option {P : : Type u) [Fintype α], Prop}
(of_equiv : β) [Fintype β] (e : α β), @P α (@Fintype.ofEquiv α β < _ > e.symm) @P β )
(h_empty : P PEmpty) (h_option : (α) [Fintype α], P α P (Option α)) : Type u)
[h_fintype : Fintype α] : P α := by
obtain p :=
let f_empty := fun i => by convert h_empty
let h_option : : Type u} [Fintype α] [DecidableEq α],
(∀ (h : Fintype α), P α) (h : Fintype (Option α)), P (Option α) := by
rintro α hα - Pα hα’
convert h_option α (Pα _)
@truncRecEmptyOption (fun α => h, @P α h) (@fun α β e hα hβ => @of_equiv α β hβ e (hα _))
f_empty h_option α _ (Classical.decEq α)
exact p _
end Fintype
/– An induction principle for finite types, analogous to Nat.rec‘. It effectively says
that every Fintype is either Empty or Option α‘, up to an Equiv‘. -/
theorem Finite.induction_empty_option {P : Type u Prop} (of_equiv : β}, α β P α P β)
(h_empty : P PEmpty) (h_option : {α} [Fintype α], P α P (Option α)) : Type u)
[Finite α] : P α := by
cases nonempty_fintype α
refine Fintype.induction_empty_option ?_ ?_ ?_ α
exacts [fun α β _ => of_equiv, h_empty, @h_option]
Instructions:
- Task 1: Modernizing a Recursion Principle for Finite Types
Consider the recursion principle for finite types that states: "To define a function or prove a property for all finite types, it suffices to handle the empty set case, show how to extend from a finite set to that set plus one extra element, and respect type equivalences."
This principle is implemented using induction on the cardinality of finite types. Refactor the implementation by replacing the manually defined recursive function with Leans standard induction tactic while preserving the mathematical content. Specifically:
1. Remove the manual recursion approach that uses the where syntax and explicit pattern matching.
2. Introduce the natural number parameter and apply induction directly.
3. In the inductive step, use the induction hypothesis provided by the tactic rather than making an explicit recursive call.
4. Ensure that all mathematical reasoning about cardinality equivalences remains intact.
The refactoring should modernize the code while maintaining the same mathematical structure and reasoning.
Generation Response of DeepSeek-R1:
<think>
Okay, I need to tackle this task of refactoring the recursion principle for finite types in the given Lean4 code. The main goal is to replace the manual recursion approach with Leans standard induction tactic while preserving the mathematical content. Let me try to break this down step by step.
First, I need to understand the existing code. The truncRecEmptyOption function uses a manual recursion approach with ind defined using the where syntax and explicit pattern matching. The comments mention that this was done to ensure computability, but the task now wants to switch to using Leans induction tactic.
The key points from the task are:
1. Remove manual recursion (‘where syntax and explicit pattern matching).
2. Introduce a natural number parameter and apply induction directly.
3. Use the induction hypothesis provided by the tactic instead of explicit recursive calls.
4. Keep the mathematical reasoning intact, especially around cardinality equivalences.
So, looking at the truncRecEmptyOption function, the manual recursion is handled by the ind function inside the apply ind where block. The ind function is defined with Nat.zero and Nat.succ n cases.
To refactor this, I need to replace this manual recursion with a standard induction tactic. That means structuring the proof so that Leans induction tactic can handle the recursion based on the natural number parameter representing the cardinality.
The suffices block is setting up the proof by showing that for all natural numbers n‘, the property P holds for ULift (Fin n)‘. Then, it binds the result of this to the equivalence between ‘α‘ and Fin (Fintype.card α)‘.
The apply ind where part is the manual recursion that needs to be replaced. Instead, I should introduce a natural number parameter (probably the cardinality) and use induction on that.
In the Nat.zero case, the current code uses truncEquivOfCardEq to get an equivalence between PEmpty and ULift (Fin 0)‘, then applies of_equiv with h_empty‘. Similarly, in the Nat.succ n case, it shows the equivalence between Option (ULift (Fin n))‘ and ULift (Fin n.succ)‘ and uses h_option with the induction hypothesis from ind n‘.
So, to use the induction tactic, Ill need to structure the proof such that:
- The induction is on the natural number n (cardinality).
- In the base case (‘n = 0‘), handle the empty set scenario.
- In the inductive step (‘n.succ‘), use the induction hypothesis to build the proof for n.succ‘.
The induction tactic should provide the induction hypothesis automatically, so I dont need to make explicit recursive calls.
I also need to ensure that all the equivalences and mathematical reasoning about cardinality are preserved. The truncEquivOfCardEq and Equiv.ulift are crucial here for maintaining those properties.
Let me sketch out the changes:
1. Remove the apply ind where block and the ind function.
2. Introduce a natural number parameter for induction, likely using Fintype.card α‘.
3. Use induction on this natural number.
4. In each case, use the existing logic with truncEquivOfCardEq and Equiv.ulift to handle the equivalences.
5. Ensure that the h_empty and h_option cases are properly applied with the induction hypothesis.
Possible challenges:
- Ensuring that the induction hypothesis is correctly applied in the inductive step.
- Maintaining the correct use of truncEquivOfCardEq and Equiv.ulift when restructuring the code.
- Making sure that all the required instances and typeclass parameters are correctly inferred or provided.
Testing the changes would involve checking that the refactored code compiles and that the mathematical content remains unchanged. Since the task specifies preserving the mathematical structure, its crucial that the equivalences and cardinality reasoning are not altered.
Another thing to note is the @[elab_as_elim]‘ attribute on induction_empty_option‘, which suggests that this theorem is intended to be used as an elimination principle. The refactoring should ensure that this attribute is still appropriate and that the theorems usability remains intact.
In summary, the main steps are:
- Replace manual recursion with induction tactic.
- Structure the proof around a natural number parameter for induction.
- Preserve all equivalences and cardinality reasoning.
- Ensure that the induction hypothesis is correctly utilized in the inductive step.
By carefully following these steps and checking against the original codes logic, I should be able to produce the required unified diff patch.
</think>
To address the task requirements, well refactor the truncRecEmptyOption function by replacing manual recursion with Leans induction tactic while preserving the mathematical content. Heres the step-by-step explanation and the corresponding diff:
**Step-by-Step Explanation:**
1. **Introduce Natural Number Parameter:**
- The original code manually recurses on the cardinality of the finite type. Well instead use a natural number parameter n and perform induction on it.
2. **Apply Induction Tactic:**
- Replace the manual ind function with induction n to leverage Leans built-in induction tactic. This simplifies the code and uses the induction hypothesis provided by the tactic.
3. **Base Case (‘n = 0‘):**
- Handle the empty set case by using truncEquivOfCardEq to establish the equivalence between PEmpty and ULift (Fin 0)‘, then apply h_empty via of_equiv‘.
4. **Inductive Step (‘n + 1‘):**
- For the successor case, use truncEquivOfCardEq to equate Option (ULift (Fin n))‘ with ULift (Fin n.succ)‘.
- Apply the induction hypothesis ih (from the induction tactic) and h_option to build the proof for n + 1‘.
5. **Preserve Equivalences:**
- Maintain the use of Equiv.ulift and truncEquivOfCardEq to ensure the cardinality reasoning remains intact.
6. **Cleanup and Simplification:**
- Remove the now-unnecessary ind function and associated comments.
- Adjust variable bindings and scopes to match the new structure.
**Unified Diff Patch:**
“‘diff
diff git a/Mathlib/Data/Fintype/Option.lean b/Mathlib/Data/Fintype/Option.lean
index 1234567..abcdefg 100644
a/Mathlib/Data/Fintype/Option.lean
$+$ b/Mathlib/Data/Fintype/Option.lean
@@ -52,29 +52,23 @@ def truncRecEmptyOption {P : Type u Sort v} (of_equiv : β}, α
(h_empty : P PEmpty) (h_option : {α} [Fintype α] [DecidableEq α], P α P (Option α))
: Type u) [Fintype α] [DecidableEq α] : Trunc (P α) := by
suffices n : ℕ, Trunc (P (ULift <| Fin n)) by
- apply Trunc.bind (this (Fintype.card α))
- intro h
- apply Trunc.map _ (Fintype.truncEquivFin α)
- intro e
- exact of_equiv (Equiv.ulift.trans e.symm) h
- apply ind where
- Porting note: do a manual recursion, instead of induction tactic,
- to ensure the result is computable
- /– Internal induction hypothesis -/
- ind : n : ℕ, Trunc (P (ULift <| Fin n))
- | Nat.zero => by
- have : card PEmpty = card (ULift (Fin 0)) := by simp only [card_fin, card_pempty,
- card_ulift]
- apply Trunc.bind (truncEquivOfCardEq this)
- intro e
- apply Trunc.mk
- exact of_equiv e h_empty
- | Nat.succ n => by
- have : card (Option (ULift (Fin n))) = card (ULift (Fin n.succ)) := by
- simp only [card_fin, card_option, card_ulift]
- apply Trunc.bind (truncEquivOfCardEq this)
- intro e
- apply Trunc.map _ (ind n)
- intro ih
- exact of_equiv e (h_option ih)
+ exact Trunc.bind (this (Fintype.card α)) fun h =>
+ Trunc.map (fun e => of_equiv (Equiv.ulift.trans e.symm) h) (Fintype.truncEquivFin α)
+ intro n
+ induction n with
+ | zero =>
+ have : card PEmpty = card (ULift (Fin 0)) := by simp only [card_fin, card_pempty, card_ulift]
+ apply Trunc.bind (truncEquivOfCardEq this)
+ intro e
+ apply Trunc.mk
+ exact of_equiv e h_empty
+ | succ n ih =>
+ have : card (Option (ULift (Fin n))) = card (ULift (Fin n.succ)) := by
+ simp only [card_fin, card_option, card_ulift]
+ apply Trunc.bind (truncEquivOfCardEq this)
+ intro e
+ apply Trunc.map _ ih
+ intro ih
+ exact of_equiv e (h_option ih)
Porting note: due to instance inference issues in SetTheory.Cardinal.Basic
I had to explicitly name h_fintype in order to access it manually.
**Explanation of the Diff:**
- **Removed Manual Recursion (‘ind function):**
- The original codes apply ind where block and manually defined ind function are deleted.
- **Introduced induction Tactic:**
- The induction n with block replaces manual recursion, using Leans built-in induction.
- **Base Case (‘zero‘):**
- Handles n = 0‘ by establishing equivalence between PEmpty and ULift (Fin 0)‘, then applying h_empty‘.
- **Inductive Step (‘succ n ih‘):**
- Uses ih (induction hypothesis) provided by the induction tactic.
- Establishes equivalence between Option (ULift (Fin n))‘ and ULift (Fin n.succ)‘.
- Applies h_option with ih to build the step case.
- **Preserved Mathematical Content:**
- All cardinality equivalence checks via truncEquivOfCardEq and Equiv.ulift remain intact.
- The structure of the proof aligns with the original mathematical reasoning.
This refactoring modernizes the code by leveraging Leans induction tactic, improving readability and maintainability while preserving the original mathematical structure.
Judgement Response of Claude Sonnet 3.7 (thinking):
<think>
Lets analyze the proposed implementation to determine if it fulfills the task requirements. First, lets understand the original code and then evaluate the changes.
## Original Code Understanding
The original code defines a recursion principle for finite types called truncRecEmptyOption‘. This function takes:
1. A way to transport through equivalences (‘of_equiv‘)
2. A base case for the empty set (‘h_empty‘)
3. An inductive step for adding one element via Option (‘h_option‘)
4. A type ‘α‘ with Fintype and DecidableEq instances
The implementation works by:
1. Reducing the problem to showing the principle for Fin n types (standard finite types of size n)
2. Using a manually defined recursive function ind to prove this for all natural numbers
3. For the base case (n=0), it establishes an equivalence with the empty type
4. For the inductive case, it uses the recursive hypothesis and establishes an equivalence between Option (Fin n)‘ and Fin (n+1)‘
The manual recursion is done using the where clause with explicit pattern matching on natural numbers.
## Task Requirements
The task asks to modernize this recursion principle by:
1. Removing the manual recursion approach with where syntax
2. Using Leans standard induction tactic
3. Using the induction hypothesis provided by the tactic
4. Preserving all mathematical reasoning
## Changes in the Proposed Implementation
Lets examine the changes:
1. **Removed manual recursion**: The implementation removed the ind function defined with the where clause and explicit pattern matching.
2. **Introduced induction tactic**: The implementation now uses induction n with to directly perform induction on the natural number parameter.
3. **Using induction hypothesis**: Instead of making an explicit recursive call via ind n‘, the implementation now uses the induction hypothesis ih provided by the tactic.
4. **Mathematical reasoning preserved**: The mathematical content about cardinality equivalences remains intact. The same logic is used in both the base case and the inductive step:
- In the base case, we establish an equivalence between PEmpty and ULift (Fin 0)‘
- In the inductive step, we establish an equivalence between Option (ULift (Fin n))‘ and ULift (Fin n.succ)‘
5. **Code cleanup**: The implementation also simplifies some nested Trunc operations by using more concise expressions:
“‘lean
exact Trunc.bind (this (Fintype.card α)) fun h =>
Trunc.map (fun e => of_equiv (Equiv.ulift.trans e.symm) h) (Fintype.truncEquivFin α)
This replaces the more verbose:
“‘lean
apply Trunc.bind (this (Fintype.card α))
intro h
apply Trunc.map _ (Fintype.truncEquivFin α)
intro e
exact of_equiv (Equiv.ulift.trans e.symm) h
Overall, the changes maintain the same mathematical structure and reasoning while modernizing the approach to use Leans built-in induction tactic rather than a manually defined recursive function.
## Evaluation
Lets assess whether the implementation correctly fulfills the task requirements:
1. **Removal of manual recursion**: The manual recursion using where and explicit pattern matching has been removed.
2. **Introduction of induction tactic**: The implementation now uses induction n with to perform induction on the natural number parameter.
3. **Use of induction hypothesis**: The implementation now uses the induction hypothesis ih provided by the tactic rather than making an explicit recursive call.
4. **Preservation of mathematical reasoning**: All mathematical reasoning about cardinality equivalences remains intact. The same logic is used in both the base case and the inductive step.
The implementation properly fulfills all the task requirements. The code has been modernized while maintaining the same mathematical structure and reasoning. The use of Leans induction tactic makes the code cleaner and more in line with modern Lean practices.
</think>
# Comprehensive Analysis of Lean4 Code Patch
## Task Understanding
The task requires modernizing a recursion principle for finite types by replacing a manual recursion approach with Leans standard induction tactic, while preserving the mathematical content. The goal is to update the code style while maintaining the same mathematical reasoning.
## Original Code Analysis
The original code implements truncRecEmptyOption‘, a recursion principle for finite types. It uses a manual recursive function definition with where syntax and explicit pattern matching on natural numbers:
“‘lean
apply ind where
ind : n : ℕ, Trunc (P (ULift <| Fin n))
| Nat.zero =>
| Nat.succ n =>
This function establishes that to define a property for all finite types, its sufficient to:
1. Handle the empty set case
2. Show how to extend from a finite set to that set plus one element
3. Respect type equivalences
The approach works by induction on the cardinality of the finite type, using equivalences between PEmpty and Fin 0‘, and between Option (Fin n)‘ and Fin (n+1)‘.
## Patch Examination
The patch makes the following changes:
1. Removes the manual recursion using where syntax and explicit pattern matching
2. Introduces direct parameter n and applies the induction tactic with induction n with
3. Uses the induction hypothesis ih provided by the tactic rather than making explicit recursive calls
4. Preserves all mathematical reasoning about cardinality equivalences
5. Streamlines some code by combining nested Trunc operations into more concise expressions
For example, this verbose nested block:
“‘lean
apply Trunc.bind (this (Fintype.card α))
intro h
apply Trunc.map _ (Fintype.truncEquivFin α)
intro e
exact of_equiv (Equiv.ulift.trans e.symm) h
Is replaced with a more concise:
“‘lean
exact Trunc.bind (this (Fintype.card α)) fun h =>
Trunc.map (fun e => of_equiv (Equiv.ulift.trans e.symm) h) (Fintype.truncEquivFin α)
## Requirement Fulfillment Analysis
The implementation successfully meets all task requirements:
1. The manual recursion approach using where syntax has been eliminated
2. The natural number parameter is introduced directly and induction is applied with the standard tactic
3. The inductive step uses the tactic-provided induction hypothesis (‘ih‘) rather than an explicit recursive call
4. All mathematical reasoning about cardinality equivalences remains intact
## Implementation Quality Analysis
The modernized implementation is:
- More idiomatic to contemporary Lean style by using built-in tactics
- Clearer in structure, as the induction flow is more explicit
- More maintainable, as it follows standard patterns that Lean developers will be familiar with
- Functionally equivalent to the original, preserving all mathematical content
- Slightly more concise due to streamlined expression of Trunc operations
## Overall Judgement
“‘json
"TaskEvaluations": {
"Task 1": "Excellent"
"FinalOverallGrade": "Excellent"